# Calling a Spade a Heart: Gaslighting Multimodal Large Language Models via Negation

### v3

@@ 2.2 Evaluation Pipeline
2.3 Gaslighting Benchmark Construction
To comprehensively evaluate the impact of negation on MLLMs, we introduce the first gaslighting benchmark, GaslightingBench, designed to ensure broad coverage, category balance, and the ability to expose MLLMs’ vulnerabilities to negation arguments. GaslightingBench exclusively uses Multiple-Choice Questions (MCQ) to facilitate structured evaluation, as MLLMs generally provide more consistent responses in MCQs than in open-ended formats, while MCQs remain more complex than binary Yes/No questions. Our benchmark is constructed through a two-step process. We first select representative questions and images. MCQs are extracted from existing datasets following three key principles: (1) We carefully review the categories across all the existing benchmarks and select a balanced representation of general categories. (2) To maintain a sufficient number of samples per category, we manually merge semantically similar categories. For example, in MMBench,“Attibute_comparison" and “Attribute_recognition" are combined under “Attribute" to ensure broader coverage. (3) For categories present in multiple datasets, we incorporate samples from different sources to ensure representativeness. For instance, our “Math" category consists of samples from both AI4MATH and MMMU-Math. Second, we employ the same approach detailed in section 2.2 to generate negation arguments for all the selected questions. Ultimately, as shown in Figure 4, GaslightingBench comprises 20 categories and 1,287 samples, covering a wide range of topics, such as Business, Medicine, Image Emotion, and Counting. Figure 5 shows a few examples from different categories in the GaslightingBench.
@@ 3 Result Analysis
MME fu2023mme MMMU yue2023mmmu MMMUPro yue2024mmmu-MMMUPro AI2Diagram kembhavi2016diagram MathVista lu2023mathvista ChartQA masry2022chartqa PoPE li2023evaluating-Pope MMBench liu2025mmbench
@@
Model LLaVA-1.6-7B liu2024llavanext-llava1.6 Qwen2-VL-7B-Instruct Qwen2VL-7B LLaVA-NeXT-8B li2024llavanext-strong Qwen2-VL-72B-Instruct Qwen2VL-7B Gemini-1.5-flash team2024gemini1.5 GPT-4o hurst2024gpt Claude-3.5-Sonnet anthropic2024claude
@@
Table 2: Results of MLLMs in our GaslightingBench, comparing each model’s performance before (i.e., initial answers) and after negation. The performance drop is highlighted in red.
Model LLaVA-1.6-7B liu2024llavanext-llava1.6 LLaVA-NeXT-8B li2024llavanext-strong Qwen2-VL-7B-Instruct Qwen2VL-7B Gemini-1.5-flash team2024gemini1.5
@@
3.1 Performance Comparison
Open-Source vs. Proprietary Models. Proprietary models generally showed better resilience. For instance, Gemini-1.5-flash has the lowest average accuracy drop 20.53%, maintaining 44.66% accuracy after negation, whereas open-source models like Qwen2-VL-72B-Instruct suffered dramatic declines, with a 46.02% drop and only manages to obtain 27.27% accuracy on average. This highlights the need for improved adversarial training in open-source MLLMs to deal with the negation arguments in conversation. In addition, proprietary models like GPT-4o do not consistently outperform open-source models across all datasets. On tasks involving true/false or multiple-choice questions, such as those in MME and AI2Diagram, GPT-4o often fails to provide a definitive answer, frequently responding with “unknown." For instance, in MME, GPT-4o incorrectly answered nearly 150 out of 400 questions related to artwork, many of which were labeled as "unknown." Furthermore, GPT-4o struggles with datasets like MathVista and ChartQA, where large open-source models like Qwen2-VL-72B outperformed it.
Result in GaslightingBench. Table 2 presents the evaluation results of MLLMs on GaslightingBench, highlighting their accuracy before and after the introduction of negation arguments. Similarly, across all models, significant performance degradation is observed, reaffirming the susceptibility of MLLMs to negation. As shown in Figure 6, different categories show varying degrees of susceptibility to negation arguments. Notably, “Social Relation" and “Image Emotion" experience the most severe drops, with accuracy declining by over 50%, indicating their high sensitivity to negation arguments. In contrast, “Geography" and “Chemistry" demonstrate comparatively lower declines but still suffer more than a 25% accuracy reduction, emphasizing that no category remains unaffected. Similar observations also appear in other existing datasets, for instance, MMMU dataset, which contains 30 subjects such as “accounting" and “Art”, the largest performance drop is observed in “Art Theory", while “Music" experiences the smallest decline. Similarly, in the MME dataset, “Counting" shows the most significant drop, whereas “Numerical Calculation" is the least affected. These results highlight the uneven but widespread impact of negation arguments across different domains.
3.2 Discussion: Effect of Negation Argument
@@
To investigate the impact of different negation strategies, we conduct experiments using two distinct negation approaches in the MMMU dataset: (1) Direct Negation Using Options. This strategy refers to directly negating the initial answer by stating the correct option explicitly, such as “The correct answer is C." (2) Negation with Descriptive Content of Option represents that we negate the initial answer by providing a description of an incorrect option, such as “… is Explicit Themes (the content of option C)." As shown in Table 3, all the MLLMs show higher accuracy for the second strategy after negation. The result suggests that providing additional context can help models better retain reasoning integrity when challenged with negation arguments. Since descriptive content more effectively demonstrates the capabilities of MLLMs compared to direct options, we adopt the second strategy for all experiments. Furthermore, as shown in Figure 8, the choice of negation arguments significantly affects MLLMs’ performance. For instance, when the negation argument includes “ten" or “eight," only two models maintain the correct response of “four," whereas using “one" as the negation argument allows six models to retain the correct answer. This highlights the varying sensitivity of MLLMs to different negation formulations. Additionally, we acknowledge that negation arguments can take numerous other forms, such as employing alternative phrasing methods. Exploring these variations will be our future work to comprehensively assess model vulnerabilities and enhance the robustness of MLLMs under diverse negation conditions.
@@
Figure 8: Example shows that different options in negation arguments (e.g., “ten," “one," and “eight") affect MLLMs’ ability to maintain their initial correct response.
4 Related Work
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) yin2023survey represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities. The introduction of contrastive learning techniques in vision-language models such as CLIP radford2021learning enables cross-modal understanding through training on extensive datasets. Building upon these foundations, modern MLLMs like GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning. Open-source counterparts, such as LLaVA li2024llavanext-strong and Qwen2-VL Qwen2VL-7B , have democratized access to advanced multimodal technologies, fostering innovation within the research community. The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions. Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct. Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications. We leave for future work the exploration of advanced reasoning foundation models, such as OpenAI’s O1 and O3, which are specifically designed to improve multi-step reasoning and tool use across modalities.
Negation Understanding. Negation, defined as the contradiction or denial of something, is a fundamental aspect of language croft1991evolution ; pea1978development . In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions kassner2019negated ; ettinger2020bert . To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets hosseini2021understanding . This line of research has advanced further with studies like truong2023language , which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation. These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts. Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding wang2023can . Recent work zhao2025aligning addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments. Building on the insights of wang2023can ; zhao2025aligning , this paper extends the investigation into the impact of opposing arguments in MLLMs.
In multimodal learning, negation has received comparatively less attention than in NLP. Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) yuksekgonul2022and ; singh2024learn ; wang2023clipn ; alhamoud2025vision . For instance, recent work alhamoud2025vision highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance. In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
5 Conclusion
This paper has revealed critical vulnerabilities in MLLMs when exposed to negation arguments in conversation when the initial answers are correct. While these models demonstrate remarkable advancements in multimodal understanding and reasoning, their susceptibility to misleading prompts highlights fundamental weaknesses in ensuring the trustworthiness of their outputs. Through comprehensive evaluation across diverse benchmarks, we emphasize the importance of addressing these challenges to enhance the reliability and robustness of MLLMs. This paper also provides valuable insights for future research aimed at developing more trustworthy AI systems capable of maintaining logical consistency under adversarial negation conditions. We encourage further exploration of advanced training techniques and evaluation frameworks to mitigate the risks posed by such adversarial inputs, ultimately advancing the state of multimodal AI toward more reliable and trustworthy real-world applications.

### v4

@@ 4.1 Performance Comparison
MME MMMU MMMUPro AI2Diagram MathVista ChartQA PoPE MMBench
@@
[…]
Table 2: Results of MLLMs in our GaslightingBench, comparing each model’s performance before (i.e., initial answers) and after gaslighting negation attack. Gemini-2.5-Pro (highlighted in gray) is evaluated in thinking mode with multi-step reasoning, whereas all other models are non-reasoning. The performance drop is highlighted in red.
Open-Source vs. Proprietary Models. Proprietary models generally showed better resilience. For instance, Gemini-1.5-flash has the lowest average accuracy drop 20.53%, maintaining 44.66% accuracy after negation, whereas open-source models like Qwen2-VL-72B-Instruct suffered dramatic declines, with a 46.02% drop and only manages to obtain 27.27% accuracy on average. This highlights the need for improved adversarial training in open-source MLLMs to deal with the negation arguments in conversation. In addition, proprietary models like GPT-4o do not consistently outperform open-source models across all datasets. On tasks involving true/false or multiple-choice questions, such as those in MME and AI2Diagram, GPT-4o often fails to provide a definitive answer, frequently responding with “unknown.” For instance, in MME, GPT-4o incorrectly answered nearly 150 out of 400 questions related to artwork, many of which were labeled as ”unknown.” Furthermore, GPT-4o struggles with datasets like MathVista and ChartQA, where large open-source models like Qwen2-VL-72B outperformed it.
Result in GaslightingBench. Table 2 presents the evaluation results of MLLMs on GaslightingBench, highlighting their accuracy before and after the introduction of negation arguments. Similarly, across all models, significant performance degradation is observed, reaffirming the susceptibility of MLLMs to gaslighting negation attack. While our main analysis focuses on non-reasoning models, we also evaluate an advanced reasoning-oriented MLLM to test the generality of this vulnerability. Specifically, we examine Gemini-2.5-Pro, currently first-ranked MLLM on the LMArena leaderboard  [^1] https://lmarena.ai/leaderboard/vision and equipped with a dedicated “thinking” mode for multi-step reasoning. Gemini-2.5-Pro achieves strong performance before negation (87.74) and shows relatively smaller degradation compared to the other non-reasoning models. However, despite its chain-of-thought reasoning, the model still exhibits notable vulnerability under gaslighting negation, with accuracy dropping to 70.86. These results reinforce our key finding: even the most capable MLLMs struggle to maintain consistent reasoning in the face of gaslighting negation, highlighting a fundamental limitation across models.
As shown in Figure 6, different categories show varying degrees of susceptibility to negation arguments. Notably, “Social Relation” and “Image Emotion” experience the most severe drops, with accuracy declining by over 50%, indicating their high sensitivity to negation arguments. In contrast, “Geography” and “Chemistry” demonstrate comparatively lower declines but still suffer more than a 25% accuracy reduction, emphasizing that no category remains unaffected. Similar observations also appear in other existing datasets, for instance, MMMU dataset, which contains 30 subjects such as “accounting” and “Art”, the largest performance drop is observed in “Art Theory”, while “Music” experiences the smallest decline. Similarly, in the MME dataset, “Counting” shows the most significant drop, whereas “Numerical Calculation” is the least affected. These results highlight the uneven but widespread impact of negation arguments across different domains.
4.2 Discussion
Effect of negation types. Negation is linguistically diverse and extends beyond the neutral forms introduced in Section 3.2. To examine this, we incorporated two additional variants into GaslightingBench: (i) anger-style negation, where the user conveys emotionally charged disbelief (e.g., “I can’t believe you made such a basic mistake!”), and (ii) authority-style negation, where the user appeals to an external authority (e.g., “The professor said your answer is incorrect.”). As shown in Table 3, we observed a larger performance decline compared to neutral negation. This indicates that emotionally or authoritatively framed challenges can further erode model reliability, likely by amplifying the model’s deference to perceived user authority.
Model LLaVA-1.6-7B Qwen2-VL-7B LLaVA-NeXT-8B Qwen2-VL-72B Gemini-1.5-flash GPT-4o Claude-3.5-Sonnet
@@ 4.2 Discussion
[…]
Table 3: Performance of MLLMs on GaslightingBench under different negation types.
Model confidence under gaslighting negation attacks. To better understand the internal behavior of MLLMs under gaslighting negation attacks, as shown in Table 4, we conduct a confidence-based analysis using model-reported probability scores for Gemini-1.5-flash and Qwen-2-VL-7B on both GaslightingBench and MMMU. We group predictions by their correctness and whether they occurred before or after negation, then compute the average confidence scores. On the one hand, for both models, confidence scores are generally higher for correct predictions than for incorrect ones, especially after negation, indicating some degree of internal calibration. On the other hand, we observe confidence drop in incorrect answers after negation, particularly in Qwen-2-VL-7B (from 90.6 to 74.4 on GaslightingBench), suggesting the model becomes less confident when misled. However, confidence in incorrect responses remains relatively high, especially in Gemini-1.5-flash, which maintains 91.9 average confidence even when wrong after negation, indicating a risk of confident hallucination. These findings indicate that while models may exhibit partial uncertainty when manipulated, they can still produce incorrect yet high-confidence outputs. This suggests the need for calibrated uncertainty modeling in MLLMs under adversarial dialogue settings.
Model Gemini-1.5-flash Qwen-2-VL-7B
Negation Before After Before After
correct incorrect correct incorrect correct incorrect correct incorrect
GaslightingBench 95.9 91.5 92.9 87.2 90.2 90.6 81.2 74.4
MMMU 93.6 91.0 94.8 91.9 90.5 91.0 84.5 87.2
Table 4: Average confidence scores of Gemini-1.5-flash and Qwen2-VL-7B on GaslightingBench and MMMU, grouped by correctness before and after gaslighting negation.
Why are MLLMs prone to gaslighting negation attacks? Our results suggest that the vulnerability of MLLMs to gaslighting-style negation stems primarily from over-alignment with human feedback in multimodal reasoning tasks. Many state-of-the-art MLLMs are trained using human preference optimization techniques, such as instruction tuning and reinforcement learning from human feedback (RLHF). While these techniques improve model helpfulness and cooperation, they also introduce a bias toward agreeing with user input, especially in conversational contexts. This can result in over-deference, where models revise initially correct answers simply in response to user disagreement, regardless of the factual correctness of the user’s claim. This behavior parallels the phenomenon of sycophancy observed in LLMs Sharma et al. (2024) and appears to extend into the multimodal setting as well. Our category-level analysis in Figure 6 further shows that the performance degradation is especially pronounced in subjective or socially nuanced categories (e.g., “Social Relation”, “Image Emotion”), whereas more objective domains like “Geography” show comparatively smaller drops. This suggests that in addition to alignment bias, task uncertainty and subjective ambiguity also play a role in model susceptibility. Our primary goal in this paper is to establish a rigorous and extensible framework for evaluating and benchmarking MLLM robustness under gaslighting negation attacks. Mitigating this issue will likely require more fine-grained alignment strategies, such as distinguishing faithful correction from invalid contradiction, and improving calibration of confidence in multimodal predictions Zhao et al. (2025). We leave these directions for future work.
5 Conclusion
This paper has revealed critical vulnerabilities in MLLMs when exposed to gaslighting negation attacks, where correct answers are overturned by misleading prompts. Our comprehensive evaluation across diverse benchmarks demonstrates that this susceptibility is widespread, affecting both proprietary and open-source models, including those with advanced reasoning capabilities. These findings highlight a fundamental gap in model robustness and raise important concerns for the trustworthiness of MLLMs in real-world applications. By introducing GaslightingBench, we provide the first systematic benchmark for evaluating this vulnerability, offering a foundation for future work on robustness and alignment. We encourage further exploration of fine-grained training strategies, confidence calibration, and evaluation frameworks that can distinguish valid corrections from misleading contradictions. Addressing these challenges is essential for advancing multimodal AI toward systems that maintain logical consistency and reliability under adversarial conditions.
Reproducibility Statement
We will release the full code, evaluation pipeline, negations for all of datasets we used, and GaslightingBench benchmark. Proprietary models were accessed via public APIs, and open-source models were evaluated with fixed checkpoints. The prompt templates are provided in the Appendix. All datasets evaluated in this paper are publicly available.
Appendix A Appendix
A.1 Performance Comparison Across Question Types
To investigate how question formats affect MLLMs’ robustness to gaslighting negation attack, we conduct a format-level analysis using the MathVista dataset, which contains Yes/No, Multiple-Choice (MCQ), and Open-Ended question types. We evaluate two models: Qwen2-VL-7B and LLaVA-NEXT-8B. As shown in the Table 5, based on the performance before negation, the relative difficulty of question formats follows the order: Open-Ended >> MCQ >> Yes/No, which aligns with expectations given that open-ended questions involve a broader and less constrained answer space. Nevertheless, open-ended questions demonstrate greater resilience to negation compared to Yes/No and MCQ formats. For example, Qwen2-VL-7B maintains 37.61% accuracy on open-ended questions after negation, compared to 26.72% for MCQs and 40.11% for Yes/No.
Model Qwen2-VL-7B LLaVA-NEXT-8B
Question Type Open-Ended MCQs Yes/No Open-Ended MCQs Yes/No
Before Negation 50.87 57.02 89.83 25.43 34.16 61.02
After Negation 37.61 26.72 40.11 14.13 9.09 19.21
Table 5: Accuracy of Qwen2-VL-7B and LLaVA-NEXT-8B on the MathVista dataset across different question formats (Open-Ended, MCQ, Yes/No) before and after gaslighting negation.
A.2 Impact of LLM Choice on Negation Generation
Although Llama3-8B-Instruct is not among the strongest models overall, we adopt it for negation prompt generation (Section 3.2) because the task is relatively constrained: the model is given the original question, answer, and candidate options (if applicable) and asked to produce a logically negated statement. In this controlled setting, even moderately sized LLMs can generate reliable outputs. To validate quality, we regenerated all negation prompts using a flagship model, Gemini-2.5-Pro, and compared them against those from Llama3-8B-Instruct. Using Qwen3 embeddings to measure semantic similarity, we obtained an average score of 0.90, confirming that the two sets of prompts are highly consistent, with negligible differences in quality.
A.3 Scope of GaslightingBench’s Question Design
The choice to base GaslightingBench primarily on Multiple-Choice Questions (MCQs) was deliberate, as discussed in Section 3.3. MCQs offer a balance of semantic complexity, evaluation consistency, and annotation reliability. They also enable precise measurement of model behavior under negation with minimal ambiguity, making them a practical foundation for probing manipulations such as gaslighting.
That said, assessing negation vulnerability across diverse question formats is equally important. Our evaluation pipeline (Section 3.2) already supports Yes/No and Free-Form questions, which we evaluate using model confidence scores or semantic similarity. Indeed, the results in Table 1 span all eight datasets, covering a mix of MCQs, binary (Yes/No), and open-ended formats (e.g., MME and MathVista).
While the initial release of GaslightingBench focuses on MCQs, we will also provide a comprehensive set of negation prompts for all eight evaluated datasets, including non-MCQ formats. This will allow the community to extend evaluations to additional question types and investigate how gaslighting negation generalizes beyond MCQs.
Model LLaVA-1.6-7B LLaVA-NeXT-8B Qwen2-VL-7B-Instruct Gemini-1.5-flash
@@ A.3 Scope of GaslightingBench’s Question Design
A.4 Effect of Negation Argument
@@
To investigate the impact of different negation strategies, we conduct experiments using two distinct negation approaches in the MMMU dataset: (1) Direct Negation Using Options. This strategy refers to directly negating the initial answer by stating the correct option explicitly, such as “The correct answer is C.” (2) Negation with Descriptive Content of Option represents that we negate the initial answer by providing a description of an incorrect option, such as “… is Explicit Themes (the content of option C).” As shown in Table 6, all the MLLMs show higher accuracy for the second strategy after negation. The result suggests that providing additional context can help models better retain reasoning integrity when challenged with negation arguments. Since descriptive content more effectively demonstrates the capabilities of MLLMs compared to direct options, we adopt the second strategy for all experiments.
In addition, we also adapt our GaslightingBench to support initial-query negation by injecting the negation statement into the original prompt. The results in Table 2 present a comparison on LLaVA-1.6-7B. We observe that initial-query negation leads to comparable or slightly worse performance than conversational negation, suggesting that negation vulnerability is pervasive regardless of prompt structure.
Model Negation GaslightingBench
LLaVA-1.6-7B Before 59.13
After(conversational negation) 27.20
Initial-query negation 21.29
Table 7: Performance of LLaVA-1.6-7B on GaslightingBench between conversational and initial-query negation.
Furthermore, as shown in Figure 8, the choice of negation arguments significantly affects MLLMs’ performance. For instance, when the negation argument includes “ten” or “eight,” only two models maintain the correct response of “four,” whereas using “one” as the negation argument allows six models to retain the correct answer. This highlights the varying sensitivity of MLLMs to different negation formulations. Additionally, we acknowledge that negation arguments can take numerous other forms, such as employing alternative phrasing methods. Exploring these variations will be our future work to comprehensively assess model vulnerabilities and enhance the robustness of MLLMs under diverse negation conditions.
@@
Figure 8: Example shows that different options in negation arguments (e.g., “ten,” “one,” and “eight”) affect MLLMs’ ability to maintain their initial correct response.
A.5 Prompt for gaslighting negation attack evaluation
The prompt template for generation negation statements and evaluation prompt templates are presented in Figures 9 and 10, respectively.
Figure 9: Prompt templates used for generating negation statements across different question formats.
Figure 10: Evaluation prompt templates used for different MLLMs across question formats.
A.6 The Use of Large Language Models (LLMs)
In this paper, LLMs/MLLMs are used in three ways: (i) Negation prompt generation. We employ Llama3-8B-Instruct to generate neutral negation prompts, chosen for efficiency and accessibility. To ensure quality, we cross-validated with Gemini-2.5-Pro and found high semantic similarity (0.90 via Qwen3 embeddings), confirming consistency across models. (ii) Model evaluation. Proprietary systems such as GPT-4o, Claude-3.5-Sonnet and Gemini-1.5/2.5 were accessed via official APIs, while open-source baselines such as Qwen-VL and LLaVA were tested on released checkpoints. All inference settings, hyperparameters, and evaluation pipelines are documented to support reproducibility. (iii) Paper writing refinement. We used LLM assistance to polish the writing and improve readability, while all technical content, methodology, and experimental design were authored and validated by the research team.